OpenStudy (anonymous):

How do you solve by substitution and graph with the problem: Eq. 1: y=1/4x + 2 and Eq. 2: y=-4x + 36?

OpenStudy (anonymous):

You need to equal the right hand side of the first equation with the right hand side of the second equation.

OpenStudy (across):

You're told that\[y=\frac{1}{4}x+2\]and\[y=-4x+36.\]Could we then say that\[\frac{1}{4}x+2=-4x+36\]and solve for \(x\)?

OpenStudy (anonymous):

And solve for x

OpenStudy (anonymous):

And then you can find the x-intercept and y-intercept of each line.

OpenStudy (anonymous):

I tried to solve for x, but i ended up with a really complicated fraction

OpenStudy (anonymous):

make x = 0 and solve for y, to find the y-incercept

OpenStudy (anonymous):

Wha did you get?

OpenStudy (across):

What if you multiply both sides by \(4\)?\[\frac{1}{4}x+2=-4x+36,\]\[x+8=-16x+144?\]Does that help with the fraction?

OpenStudy (anonymous):

She multiply both sides of the equation by 4

OpenStudy (anonymous):

multiplied

OpenStudy (anonymous):

thank you so much!! Because i did that, i found out what x is. How would you solve by graphing?

OpenStudy (anonymous):

You can follow this steps: - substract 8 from both sides of the equation. - add 16x to both sides of the equation. - divide by 17 both sides of the equation. And I think (if i'm not wrong) you get the value of x that satisfies the problem.

OpenStudy (anonymous):

ahh.

OpenStudy (anonymous):

Do you know how to graph a linear equation?

OpenStudy (anonymous):

so you dont have to solve for y? you can solve for x to graph it?

OpenStudy (anonymous):

Once you get the value of x, you can plug in to one of your equations and solve for y.

OpenStudy (anonymous):



OpenStudy (anonymous):

Did you get the value of y?

OpenStudy (anonymous):

hold on, im still working the problem out

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

to graph, do i have to start from 36 because the y intercept is 36 for y=-4x+36

OpenStudy (anonymous):

yes if you let x = 0

OpenStudy (anonymous):

so you have the coordinate (0,36)

OpenStudy (anonymous):

You need the coordinates of another point on the line to graph it. You can get it with the x-intercept.

OpenStudy (anonymous):

let y = 0, and solve for x.

OpenStudy (anonymous):

so the x intercept would be 9 right?

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

So you have another coordinate (9,0)

OpenStudy (anonymous):

You can do a line thar passes across those points.

OpenStudy (anonymous):

and you can do the same for the other line.

OpenStudy (anonymous):

pass, sorry

OpenStudy (anonymous):

I'm not proud of my english =P

OpenStudy (anonymous):

but your awsome at math :) if that makes you feel any better

OpenStudy (anonymous):

haha thank you.

OpenStudy (anonymous):

thank u :) this really helped alot

OpenStudy (anonymous):

You're welcome =)